OpenStudy (anonymous):

Find the value of: \(\huge\color{blue}{(\frac{sec(\theta) + tan(\theta) - 1}{tan(\theta) - sec(\theta) + 1})^2}\).

OpenStudy (anonymous):

@Directrix @hartnn

OpenStudy (anonymous):

@jim_thompson5910 @campbell_st @dumbcow

OpenStudy (anonymous):

Expand then use some trigo identities to simplify it. The answer is sec(theta)+tan(theta).

OpenStudy (anonymous):

Certainly I don't want answer, I have answer choices with me, I just want the solution.. Simply tell me what formula to use here..

OpenStudy (anonymous):

oops, sorry, should be (sec(theta)+tan(theta))^2

OpenStudy (anonymous):

Never mind, don't concentrate on answer...

OpenStudy (anonymous):

\[ =\frac{\sec^2\theta+2\sec\theta(\tan\theta-1)+(\tan\theta-1)^2}{\sec^2\theta-2\sec\theta(\tan\theta+1)+(\tan\theta+1)^2}\\ =\frac{\sec^2\theta+2\sec\theta\tan\theta-2\sec\theta+\tan^2\theta-2\tan\theta+1}{\sec^2\theta-2\sec\theta\tan\theta-2\sec\theta+\tan^2\theta+2\tan\theta+1}\\ =\frac{2\sec^2\theta+2\sec\theta\tan\theta-2\sec\theta-2\tan\theta}{2\sec^2\theta-2\sec\theta\tan\theta-2\sec\theta+2\tan\theta}\\ =\frac{\sec\theta(\sec\theta+\tan\theta)-(\sec\theta+\tan\theta)}{\sec\theta(\sec\theta-\tan\theta)-(\sec\theta-\tan\theta)}\\ =\frac{\sec\theta+\tan\theta}{\sec\theta-\tan\theta} \]

OpenStudy (anonymous):

sec(theta)=1/cos(theta) and tan(theta)=sin(theta)/cos(theta), first step is trying to rewrite the given expression in terms of sin(theta) and cos(theta)

OpenStudy (anonymous):

\[=\frac{1+\sin\theta}{1-\sin\theta}\]

OpenStudy (anonymous):

I am sure I might have messed up somewhere in the tex!!

OpenStudy (anonymous):

Yep you are right electrokid... Thank you so much.. My mind was not working on it earlier.

OpenStudy (anonymous):

yay!

OpenStudy (anonymous):

Let me check it now carefully, thanks to all for coming to atleast see this post..

OpenStudy (anonymous):

when a smrty with score 99 posts a question, we expect some fun :D

OpenStudy (anonymous):

Was this not fun??